# RADLADS: Rapid Attention Distillation to Linear Attention Decoders at Scale

### v3

@@ 6 Ablation Studies
Arch ablation lambada MMLU arc_c arc_e hella piqa winog RAD-RWKV6 none 0.6748 0.6572 0.5631 0.8136 0.7901 0.8025 0.7111 RAD-RWKV6 use rope 0.6740 0.6610 0.5666 0.8165 0.7886 0.7992 0.7056 RAD-RWKV6 no tokenshift 0.6707 0.6584 0.5546 0.8043 0.7874 0.8036 0.6875 RAD-RWKV6 no gate 0.6590 0.6417 0.5444 0.8077 0.7842 0.7971 0.6930 RAD-RWKV6 use groupnorm 0.6559 0.6340 0.5648 0.8186 0.7865 0.7905 0.7032
7 Training Details and Hyper-parameters
We use the DCLM (Li et al., 2024b) dataset for all three steps in the RADLADS attention conversion process.
Step Tokens LR LR decay seqlen batch size adam betas, eps 1 100M cosine 512 32 0.9, 0.95, 2 250M-700M none 512 96 0.9, 0.95, 3 100M none 16384 96 0.9, 0.95,
7.1 General
@@ 8 What Did Not Work
Initial Attention Score Alignment
We originally included an initial ”step 0”, which was similar to step 1 but used a loss based on attention scores differences. For linear RNNs of the form
@@
This generalizes Mamba2 (Dao & Gu, 2024), RWKV5, RWKV6 (Peng et al., 2024), RWKV7 (Peng et al., 2025), DeltaNet (Schlag et al., 2021), and Gated DeltaNet (Yang et al., 2025), as well as others. See Table 1 in Peng et al. (2025) for a detailed list of such RNNs.
@@ 9 Conclusions
However, RADLADS does have known limitations. Currently, each new architecture design requires meticulous testing to improve its compatibility with the RADLADS protocol. As one example, RAD-RWKV7 appears to exhibit reduced training stability at larger (32B+) parameter and layer counts, and we are actively working to design architectural and training methodology changes to remedy this. New varieties of model interactions will test these conversions in ways that we cannot predict. For example, the details of how the conversion process impacts reasoning models like our converted QwQ model is still unknown and requires further testing.
We have many ideas for future work to enhance and extend the RADLADS process. Converting to and from different architectures and testing of more varieties of datasets against each of these would help discover the principles for optimal conversion dataset design. We also believe that parts of our existing RAD-RWKV7 architecture can still be improved; balancing of state to create implicit normalization and then removal of groupnorm could further increase downstream performance.
@@ Acknowledgments
Eric Alcaide acknowledges support for this work from the Swiss State Secretariat for Education, Research and lnnovation (SERI) under contract number 23.00421.
Appendix A Author Contributions
Appendix B Architecture Detail
@@
B.1 RAD-RWKV6 Time Mixing
RAD-RWKV6 is a customized variation of RWKV6-C2 (”Finch-C2”) (Goldstein et al., 2024), featuring a Gated Linear Attention (Yang et al., 2024) kernel (so no bonus), and a sigmoid gate. The use of the state balancing technique from RWKV6-C2 enables us to remove state normalization, improving downstream performance. Removal of bonus did not harm downstream performance. We do not typically use RoPE in this formulation.
@@ B.1 RAD-RWKV6 Time Mixing
Time mixing is performed by the attention calculation:
@@
B.2 RAD-RWKV7 Time Mixing
RAD-RWKV7 is a customized variation of RWKV-7 (Peng et al., 2025), with tokenshift removed, RoPE applied, and no bonus. The removal of tokenshift speeds up training and inference, and bonus had no beneficial impact on downstream performance.
@@ B.2 RAD-RWKV7 Time Mixing
Before mixing in the time dimension, is normalized per head:
| normalized removal key | (28) |
Time mixing is performed by the attention calculation:

### v4

@@ 7 Ablation studies
Arch ablation lmbda mmlu arc_c arc_e hella piqa winog RAD-RWKV6 none 67.48 65.72 56.31 81.36 79.01 80.25 71.11 RAD-RWKV6 use rope 67.40 66.10 56.66 81.65 78.86 79.92 70.56 RAD-RWKV6 no tokenshift 67.07 65.84 55.46 80.43 78.74 80.36 68.75 RAD-RWKV6 no gate 65.90 64.17 54.44 80.77 78.42 79.71 69.30 RAD-RWKV6 use groupnorm 65.59 63.40 56.48 81.86 78.65 79.05 70.32
@@ 8 What did not work
Initial attention matrix orientation
We originally included an initial ”step 0,” which was similar to step 1 but used a loss based on attention scores differences. For linear RNNs of the form
@@
This generalizes to Mamba2 (Dao and Gu, 2024), RWKV5, RWKV6 (Peng et al., 2024), RWKV7 (Peng et al., 2025), DeltaNet (Schlag et al., 2021), and Gated DeltaNet (Yang et al., 2025), as well as others. See Table 1 in Peng et al. (2025) for a detailed list of such RNNs.
@@
GroupNorm
Many recent RNN models rely on GroupNorm or LayerNorm to normalize the heads after the attention operation instead of using a denominator as in softmax attention or the classic linear attention formula. We found that for models around 14B and higher, this began to cause serious training instabilities. After examining gradients we theorized that this is because late normalization allows unbounded growth of the inputs to that normalization, which can become catastrophic for deep models. As a result, we applied pre-scaling to keys as in RWKV-6c and a new extension that we developed for RWKV-7, allowing us to remove the normalization and denominator.
@@
Weight tying
We initially had trouble obtaining good results when distilling smaller models featuring tied embedding weights. Eventually we found a workaround, which was to allow the embeddings and language modeling head to be untied during distillation, but then discard the updated head at the end and retie the weights. We are uncertain why this worked or whether it will act consistently across architectures. In order to avoid this confounding factor we distilled only models containing untied weights for this paper.
@@
9 Distillation and student architectures
We theorize that mechanisms not present in the distilled teacher model (e.g. tokenshift/convolution) offer limited benefit to a distilled student. With longer training these mechanisms could become useful for the converted model. However, our extremely short distillation process probably prevents enough optimization from occurring for the model to learn to put them to good use. Conversely, mechanisms from the teacher (e.g. RoPE) can be important not because they are strictly necessary, but because there are not enough distillation steps for the model to learn to replace them fully. Similarly, we theorize that GQA is probably not an overall benefit in terms of downstream results with enough training, but because the Qwen teacher employed it, it provides a good starting point and guidance for the model during distillation.
@@ 10 Conclusions
However, RADLADS does have known limitations. Currently, each new architecture design requires meticulous testing to improve its compatibility with the RADLADS protocol. As one example, the GroupNorm or LayerNorm often used in place of a denominator in linear attention variants caused severe training instability at 14B+ scale during distillation and needed replacement. New varieties of model interactions will test these conversions in ways that we cannot predict.
We have many ideas for future work to enhance and extend the RADLADS process. Converting to and from different architectures and testing of more varieties of datasets against each of these would help discover the principles for optimal conversion dataset design. A future direction could be to have the teacher model generate synthetic data, to ensure that the dataset used for conversion remains in-distribution, especially for reasoning models. Another interesting direction is conversion to hybrid models like Goldstein et al. (2024), including conversion that takes into account sparsity or methods of KV Cache compression.
@@ Acknowledgments
We would like to thank TensorWave and AMD for sponsoring the initial compute used for this project to train models on MI300X hardware. Eric Alcaide acknowledges support for this work from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number 23.00421.
Appendix A Architecture detail
@@
A.1 RAD-RWKV6 sequence mixing
RAD-RWKV6 is a customized variation of RWKV6-C2 (”Finch-C2”) (Goldstein et al., 2024), featuring a Gated Linear Attention (Yang et al., 2024) kernel (so no bonus), and a sigmoid gate. The use of the state balancing technique from RWKV6-C2 enables us to remove state normalization, improving training stability and downstream performance. Removal of bonus did not harm downstream performance. We do not typically use RoPE in this formulation.
@@ A.1 RAD-RWKV6 sequence mixing
Sequence mixing is performed by the attention calculation:
@@
A.2 RAD-RWKV7 sequence mixing
RAD-RWKV7 is a customized variation of RWKV-7 (Peng et al., 2025), with tokenshift removed, RoPE applied, and no bonus. The removal of tokenshift speeds up training and inference, and bonus had no beneficial impact on downstream performance. The use of a state balancing technique from RWKV6-C2, modified to apply properly to RWKV-7, enables us to remove state normalization, improving training stability.
| decay precursor | (19) | ||||
@@ A.2 RAD-RWKV7 sequence mixing
Before mixing in the time dimension, is normalized per head to form the removal key :
Sequence mixing is performed by the attention calculation:
@@
A.3 Linear Attention sequence mixing
The Linear Attention used to train the QLinAtt-7B-Instruct model in the results section is a slight variation of traditional linear attention. It uses no function, and applies LayerNorm per head to normalize the attention results, formally stated per head as:
Finally, the heads are recombined via reshaping so that , and transformed into the output as follows:
Appendix B Additional training settings
Appendix C Absolute scores summary
Student / Teacher accuracy (%) Name lmbda mmlu Others SUPRA 69.2 / 75.8 33.1 / 62.4 69.8 / 74.4 LOLCats / Hedgehoga - 23.8 / 66.6 61.6 / 73.1 MOHAWK 50.1 / 53.4 24.2 / 41.8 65.1 / 67.2 Mamba in the Llama 43.9 / 72.2 45.2 / 63.9 65.4 / 72.9 DiJianga - 40.7 / 46.8 - ARWKV 62.2 / 69.6 62.4 / 71.7 71.3 / 73.5 Llamba 69.4 / 73.0 61.0 / 68.0 73.8 / 74.2 RADLADS (ours) 68.4 / 69.6 68.2 / 71.7 73.7 / 73.5
Appendix D Advanced and long-context benchmarks
After using the RADLADS protocol to create the RNN models featured in this paper, we became interested in trying the technique to adapt teacher models to non-RNN forms. To this end, we distilled Qwen3-8B-Base to utilize a new attention variant named Softpick attention(Zuhri et al., 2025). The results for select advanced benchmarks are shown in Tables 10 and 12.
Performance was excellent, rivaling or exceeding the original teacher model on every test. This was encouraging, since Softpick attention is quite different from normal attention, and removes all negative attention scores. However, the distillation failed to achieve the primary goal of Softpick attention, which is to avoid the creation of attention sinks. Analyzing the attention maps showed that our distilled model still contained these sinks. Despite not achieving the goals of the Softpick architecture in this case, we believe that the ability to inexpensively convert to new non-RNN forms of attention may be a useful tool for researchers.
model gsm8k(%) HumanEval(%) passkey¿=95 (ktok) QRWKV7-Qwen2.5-7B 61.4 / 85.9 50.0 / 79.9 8k / 34k+ ARWKV 53.4 / 85.9 50.0 / 79.9 1k / 34k+ Llamba-8B 56.0 / 81.0 6.1 / 56.1 1k / 128k* (*Meta’s claim) Softpick-Qwen3-8B-Base 85.7 / 86.1 81.7 / 80.5 34k+ / 34k+
These QRWKV and Softpick models were distilled using only steps 1 and 2a
GSM8K benchmark was run with max_new_tokens=512
model ctx RULER score QRWKV7-7B-Instruct 4096 39.7 QRWKV7-7B-Instruct 8192 26.1 QRWKV7-7B-Instruct 16384 16.5 teacher: Qwen2.5-7B-Instruct 16384 69.7
Model Max allowed response length Accuracy QRWKV7-Qwen3-8B 1024 25.9 teacher: Qwen3-8B 1024 41.4 QRWKV7-Qwen3-8B 2048 27.6 teacher: Qwen3-8B 2048 59.9 Qwen3-Softpick-8B-Base 1024 53.0 teacher: Qwen3-8B-Base 1024 53.1
Appendix E Inference speed
In order to demonstrate the improvement in inference speed achievable by converting from Qwen to QRWKV7 via the RADLADS process, we test various context lengths using an optimized vLLM instance on 1x H100 in bfloat16, with 32 simultaneous requests. We find that with as few as 8k tokens in and 256 tokens out, there is a significant speed increase. This increase continues to widen as the number of output tokens increases. RAD-RWKV embeddings and MLP layers are the same shapes and use the same computation as the teacher model. Therefore, all speed improvement comes from the attention replacement, which uses a fixed size state and much lower memory bandwidth.
Tokens in Tokens out QRWKV7-32B Qwen3-32B ratio 8k 256 86.5 sec 139.5 sec 1.61x 7k 1k 148.7 sec 411.7 sec 2.77x 6k 2k 229.6 sec 783.0 sec 3.41x